3x - 1 - x = 3x + 5, -4(x + 7) = -16, 5(x + 3) = 7(x - 5)
Answer by daffy_duck(3)

1) 3x-1-x=3x+5
-x=6
x=6
2) -4(x+7)=-16
-4x-28=-16
-4x= 12
x=-3
3) 5(x+3)=7(x-5)
5x+15=7x-35
50=2x
25=x